Please help me to figure out how -2p+8=3 is y=5/2
Answer by richwmiller(17219)

It isn't.
p=5/2
not y=5/2
-2p+8=3
subtract 8 from both sides
-2p=-5
divide by -2
p=-5/-2
p=5/2